# VerifiableFL: Verifiable Claims for Federated Learning using Exclaves

### v3

@@ Abstract.
In federated learning (FL), data providers jointly train a machine learning model without sharing their training data. This makes it challenging to provide verifiable claims about properties of the final trained FL model, e.g., related to the employed training data, the used data sanitization, or the correct training algorithm—a malicious data provider can simply deviate from the correct training protocol without being detected. While prior FL training systems have explored the use of trusted execution environments (TEEs) to combat such attacks, existing approaches struggle to link attestation proofs from TEEs robustly and effectively with claims about the trained FL model. TEEs have also been shown to suffer from a wide range of attacks, including side-channel attacks.
We describe VerifiableFL, a system for training FL models that provides verifiable claims about trained models with the help of runtime attestation proofs. VerifiableFL generates such proofs using the new abstraction of exclaves, which are integrity-only execution environments without any secrets, thus making them immune to data leakage attacks. Whereas previous approaches only attested whole TEEs statically, i.e., at deployment time, VerifiableFL uses exclaves to attest individual data transformations during FL training. These runtime attestation proofs then form an attested dataflow graph of the entire FL model training computation. The graph can be checked by an auditor to ensure that the trained FL model satisfies its verifiable claims, such as the use of particular data sanitization by data providers or aggregation strategy by the model provider. We implement VerifiableFL by extending NVIDIA’s NVFlare FL framework to use exclaves, and show that VerifiableFL introduces less than 10% overhead compared to unprotected FL model training.
@@ 1. Introduction
As machine learning (ML) models continue to be used in new and diverse applications, there is increasing impetus towards providing information about how these models were trained. The EU Artificial Intelligence Act (EUA, 2024), which came into force in August 2025, requires providers of general-purpose ML models to be able to provide “information on the data used for training, […] including the type and provenance of data and curation methodologies (e.g. cleaning, filtering etc.), the number of data points, their scope and main characteristics; how the data was obtained and selected as well as all other measures to detect the unsuitability of data sources and methods to detect identifiable biases, where applicable”.
@@
In response to such requirements, datasheets for datasets (Gebru et al., 2021) describe the composition of training datasets, and model cards (Mitchell et al., 2019) include a section that describes model training. Providers of open-source models, such as the recently-trained Apertus (Hernández-Cano et al., 2025), publish the code and datasets used for training; providers of closed-source models make various claims about their training process and datasets: e.g., the model card for OpenAI’s GPT-5 describes how the training data was pre-processed to remove personal information (OpenAI, 2025). Such claims about model training can be used by model providers as well as any downstream users. For example, model providers may want to demonstrate to a regulator that a model has not been trained on personal data or copyrighted content. Alternatively, they may want to show that the model has been trained on certain data, such as under-represented languages, as claimed by Apertus (Hernández-Cano et al., 2025).
@@
At present, any claims about model training and the employed datasets are made without proof and thus cannot be verified. For closed-source models, it is impossible to verify claims when interacting with the model via an inference API. Even for open-source models, it is generally infeasible to verify that a given model is indeed the result of executing any provided training code on a specified dataset—doing so would essentially require repeating the training process, which would incur substantial costs, and the trained model may still differ due to the non-determinism in the training process. Today, claims about models are therefore accepted based on the reputational trust in model providers.
@@
The situation becomes even worse with federated learning (FL) (McMahan et al., 2017; Kairouz et al., 2021; Jere et al., 2020), because the training is performed by (potentially many) data providers, rather than a single model provider. Any claims about models trained using FL therefore require reputational trust in all data providers, as well as the model provider. Nevertheless, FL is an appealing approach in many settings, because it allows data providers to retain control over datasets. Its use has been proposed for applications ranging from training large predictive language models on user input (Hard et al., 2018; Ramaswamy et al., 2019), to training fraud detection models (Yang et al., 2019) while protecting transaction data of competing financial institutions.
@@
In this paper, we ask the question whether it is possible “to provide verifiable claims about how an ML model was trained using FL by multiple data providers”. We observe that this would require systems support during training to discover any deviations by parties from the agreed training protocol, which would invalidate particular claims. For example, the system would have to discover if a data provider skips a given pre-processing step of a dataset that removes personal information. There could be various reasons for such deviations, ranging from accidental  (e.g., bugs in the code) through to actively malicious (e.g., deliberately including prohibited data to poison the model (Tolpegin et al., 2020)).
@@
A strawman solution for creating an FL training system that offers verifiable claims about the execution of an FL training job would be to execute the FL training computation inside trusted execution environments (TEEs), such as Intel SGX enclaves (Arnautov et al., 2016) or AMD SEV-SNP confidential virtual machines (Sev-Snp, 2020). TEEs ensure that the integrity and confidentiality of computation is preserved by the hardware: they isolate the TEE memory used for the computation and offer attestation proofs, which contain a signed measurement of the TEE memory taken at deployment time. Indeed, prior FL training systems have employed TEEs to offer claims about trained ML models by leveraging attestation proofs about the executed computation (ope, 2025; Quoc and Fetzer, 2021; Guo et al., 2024).
@@
We argue, however, that TEEs are the wrong abstraction to create verifiable claims about model training for two reasons:
@@
(1) Focus on static attestation proofs. In current FL training systems that use TEEs (ope, 2025; Quoc and Fetzer, 2021; Guo et al., 2024), TEEs produce memory measurements once at deployment time to provide a static attestation proof that establishes that the TEE runs the expected FL system and training code. A model provider can then check each data provider’s TEE attestation e.g., when performing an initial TLS handshake (ope, 2025). Such a static attestation approach, however, makes it challenging to relate these attestations to any claims about the final trained model: verifiable claims depend on how the training process was carried out at runtime but a static attestation proof only states that the training started from a known state. For example, if a data provider skips data sanitization in a given training round to poison the trained model, this would be difficult to discern from a single static attestation proof.
(2) Over-reliance on confidentiality. TEEs attempt to guarantee confidentiality by isolating data within TEEs from external access. Unfortunately, this guarantee is not absolute e.g., it only holds under the threat model where advanced attacks such as side channels are out of scope. Recent research (Van Bulck et al., 2018; Schwarz et al., 2019; Li et al., 2022) has shown that side-channels are a real concern. However, as we show, virtually all TEE implementations and applications conveniently rely on confidentiality, even for seemingly unrelated properties such as integrity of execution and unforgeability of attestation. As a result, these applications are exposed to adversaries that can use side-channel attacks: e.g., if a distributed FL training system relies on attested TLS connections between data providers and the model provider (Guo et al., 2024), an adversary who manages to leak the TLS session key from a TEE using side channels can deviate from the correct training protocol.
To address these issues, we describe a new abstraction called an exclave, which is a hardware-based integrity-only execution environment. Exclaves preserve execution integrity and are designed to attest data transformations, such as data sanitization (Padhi et al., 2024), local training (Kairouz et al., 2021),
and model aggregation (Yin et al., 2018) carried out during FL training. By providing a set of runtime attestation proofs of data transformations, it becomes possible to relate these proofs to claims about the final trained FL model.
A distinguishing feature of exclaves is that their implementations do not rely on secrets held in software, and require that programmers design applications with no secrets in software either. In particular, exclaves do not rely on secrets for attestation, and require that programmers do not rely on primitives such as TLS for integrity. To support attestation, exclaves exclusively use hardware-based attestation keys to sign attestation statements. The attestation key is managed entirely in hardware (e.g., in a secure co-processor (Sev-Snp, 2020)), otherwise inaccessible to software. This design makes exclaves resilient to data leakage attacks, such as side-channel attacks. We show how exclaves can be realized through minor modifications to existing hardware-based TEE implementations.
Using exclaves, we describe the design and implementation of VerifiableFL, a system for executing the FL training job that provides verifiable claims about the trained FL model. VerifiableFL extends an existing FL framework, such as NVFlare (Roth et al., 2022), to execute FL tasks, which carry out data transformation as part of an FL training job, inside exclaves. VerifiableFL achieves this without requiring changes to the FL jobs or the FL framework implementation.
@@
VerifiableFL uses exclaves to generate runtime attestation proofs of all executed data transformations in the FL job, which covers the transformations performed by the data providers and the aggregations performed by the model provider in each training round. Each attestation proof, which we call an exclave data record (EDR), binds together (through hashes): (i) the data inputs to the transformation; (ii) the code performing the transformation; and (iii) the resulting data output of the transformation. The EDRs are subsequently collected by VerifiableFL and used to construct an end-to-end attested exclave dataflow graph (EDG), which provides a complete and precise representation of all data transformations that led to the final trained FL model. The EDG can be used by a third-party auditor to verify claims about the FL model, e.g., that data pre-processing and model training has been carried out faithfully by data providers across rounds and that the model provider has aggregated all model contributions by the data providers without bias.
@@
We implement a prototype of VerifiableFL using confidential computing hardware and services on the Azure cloud environment (Russinovich et al., 2021). Our experiments show that VerifiableFL can provide verifiable claims such that the FL model was trained with the correct algorithms and datasets with low overhead: the execution of the FL training job within exclaves and the generation of runtime attestation proofs increases training time for compute-intensive models such as VGG9, BERT, and GPT-2 by less than 10% compared to a baseline without exclaves.
@@ 2. Verifiable Claims in Federated Learning
Next, we provide background on training machine learning models using federated learning (§2.1) and discuss the challenges when providing verifiable claims about trained models (§2.2). After that, we introduce trusted execution environments (§2.3) and survey how such approaches and other algorithmic techniques can provide verifiable claims about trained models (§2.4).
@@ 2.1. Federated learning
In FL training, other techniques exist to ensure the protection of datasets or the quality of the trained FL model. For example, data providers may apply techniques such as differential privacy (DP) (Dwork et al., 2014) to bound the contribution of individual records in the training dataset. In DP, a data provider adds noise to the model updates in order to protect the privacy of its dataset. To ensure the quality of the trained FL model, the FL training jobs may include algorithmic approaches to detect or mitigate against poisoning attacks (Tolpegin et al., 2020), e.g., using outlier detection (Tolpegin et al., 2020) or robust aggregation (Blanchard et al., 2017; Yin et al., 2018).
2.2. Claims about trained models
There is an increasing requirement to provide information and make claims about various aspects of the training process for ML models. Prior work has presented proposals for providing such information, which are now being used in practice, such as datasheets for datasets (Gebru et al., 2021) and model cards (Mitchell et al., 2019). Similarly to their use in electronics, the idea of datasheets for datasets is to describe the composition and intended use of a training dataset. Model cards perform a similar function, but for a trained model, as shown in the example in Fig. 2. Notably, model cards should ideally include information such as which training algorithm was used, which training datasets were used (e.g., by including pointers to the datasets’ datasheets), as well as details of any pre-processing performed on the datasets before training (e.g., to check for and/or filter out certain types of content).
@@ 2.2. Claims about trained models
In fully open-source models, the exact training code and dataset is usually made publicly available. In open-weight or closed-source models, the training code and dataset may be confidential, but model providers may still make public claims about the training code, dataset, and any data pre-processing performed. For example, the card for OpenAI’s GPT-5 model claims that their data processing pipeline “includes rigorous filtering to maintain data quality and mitigate potential risks” and that they use “advanced data filtering processes to reduce personal information from training data” (OpenAI, 2025). For all types of models, information about training may need to be provided to regulators. For example, Article 53 of the EU AI Act requires model providers to provide this information, “upon request, to the AI Office and the national competent authorities” (EUA, 2024).
Making claims about training with FL, e.g., that particular datasets have been used or attack mitigations have been in place, is even more challenging. Claims about a trained FL model require trust in the model provider and all data providers who contributed to the training. There could be various reasons for any of these participants to deviate from the agreed training protocol, both unintentionally and intentionally. For example, a data provider may accidentally omit or misconfigure steps in pre-processing the training data, which could lead to undesirable data being used for training. Alternatively, to reduce costs, a data provider may promise to train on a particular dataset (e.g., patient data from a hospital) but might only train on a subset of this data, leading to a discrepancy between the claimed and actual training dataset.
It is currently challenging to verify claims about model training, even with access to the training code and dataset. For closed-source models that can only be accessed via inference APIs, there is no way to verify claims about the model training. For open-source models, one could attempt to repeat the training process, but depending on the size of the model and the length of training, this may require significant hardware resources and thus incur substantial costs. Even if this were feasible, there is no guarantee that the trained model would match the original model, due to the inherent non-determinism in the training process. Therefore, claims about model training typically have to be accepted based on the reputational trust in the model provider making the claim. Organisations would experience reputational damage if their claims about trained ML models were found to be incorrect.
@@ 2.3. Trusted execution environments
Trusted execution environments (TEEs) try to ensure both the confidentiality and integrity of computation in a verifiable manner. They isolate the computation and its data, preventing even privileged attackers, such as malicious administrators, operating systems (OS), or hypervisors, from accessing the data or tampering with execution. TEEs also support attestation, which enables external parties to verify a measurement taken over the TEE memory that includes both its code and data. TEEs, including attestation, are implemented using hardware-based mechanisms, such as memory encryption, memory access checks, and hardware signatures for attestation.
TEEs exist at different granularities: ARM TrustZone (Ngabonziza et al., 2016) provides a separate trusted world; Intel Software Guard Extensions (SGX) provides isolation at a process granularity. More recently, VM-based TEEs, such as AMD SEV-SNP (Sev-Snp, 2020), Intel TDX (Witharana et al., 2024; Cheng et al., 2023), and ARM CCA (ARM, 2024; Li et al., 2023), isolate entire virtual machines (VMs) from the hypervisor and other privileged software.
Hardware support for TEEs has also been integrated into GPUs. For example, NVIDIA’s H100 GPUs support a confidential mode with a unified TEE, spanning the CPU-based TEE and the GPU (Dhanuskodi et al., 2023; Roth et al., 2022; Hande, 2023). All sensitive GPU state, including GPU memory and configuration registers, are protected from external access by the hardware, and all communication between the CPU-based TEE and the GPU is encrypted and integrity-protected. Graphcore IPUs provide similar TEE functionality (Vaswani et al., 2023).

### v4

# VerifiableFL: Providing Verifiable Claims to Federated Learning using Exclaves

@@ Abstract.
In federated learning (FL), data providers jointly train a machine learning model without sharing their training data. This makes it challenging to provide verifiable claims about the trained FL model, e.g., related to the employed training data, any data sanitization, or the correct training algorithm—a malicious data provider can simply deviate from the correct training protocol without detection. While prior FL training systems have explored the use of trusted execution environments (TEEs) to protect the training computation, such approaches rely on the confidentiality and integrity of TEEs. The confidentiality guarantees of TEEs, however, have been shown to be vulnerable to a wide range of attacks, such as side-channel attacks.
We describe VerifiableFL, a system for training FL models that establishes verifiable claims about trained FL models with the help of fine-grained runtime attestation proofs. Since these runtime attestation proofs only require integrity protection, VerifiableFL generates them using the new abstraction of exclaves. Exclaves are integrity-only execution environments, which do not contain software-managed secrets and thus are immune to data leakage attacks. VerifiableFL uses exclaves to attest individual data transformations during FL training without relying on confidentiality guarantees. The runtime attestation proofs then form an attested dataflow graph of the entire FL model training computation. The graph is checked by an auditor to ensure that the trained FL model satisfies its claims, such as the use of data sanitization by data providers or correct aggregation by the model provider. VerifiableFL extends NVIDIA’s NVFlare FL framework to use exclaves. Since exclaves require minor yet proprietary hardware changes, we implement a variant of our design with similar performance characteristics but slightly weaker security guarantees on existing AMD SEV-SNP hardware. We show that VerifiableFL introduces less than 12% overhead compared to unprotected FL model training.
@@ 1. Introduction
As machine learning (ML) models continue to be used in new and diverse applications, there is increasing impetus towards providing information about how these models were trained. For example, the EU AI Act (EUAIAct), which came into force in August 2025, requires providers of general-purpose ML models to provide “information on the data used for training, […] including the type and provenance of data and curation methodologies (e.g. cleaning, filtering etc.), the number of data points, their scope and main characteristics; how the data was obtained and selected as well as all other measures to detect the unsuitability of data sources and methods to detect identifiable biases, where applicable”.
@@
In response to such requirements, datasheets for datasets (datasheetsdatasets) describe the composition of training datasets, and model cards (modelcard) describe model training. Providers of open-source models, such as the recently-trained Apertus (apertus), publish the code and datasets used for training; providers of closed-source models also make claims about their training process and datasets: e.g., the model card for OpenAI’s GPT-5 describes how the training data was pre-processed to remove personal information (gpt5). Such claims may be relevant to model providers as well as downstream users. For example, model providers may want to demonstrate to regulators that a model has not been trained on personal data or copyrighted content; alternatively, they may want to show that the model has been trained on certain data, such as under-represented languages, as claimed by Apertus (apertus).
@@
At present, such claims about model training and datasets are made without proof and thus cannot be verified. For closed-source models, it is impossible to verify claims when interacting with the model via an inference API. Even for open-source models, it is generally infeasible to verify that a given model is indeed the result of executing the provided training code on a specified dataset—doing so would essentially require repeating the training process, which would incur substantial costs, and the trained model may still differ due to non-determinism. Today, model claims are therefore accepted based on the reputational trust in model providers.
@@
In federated learning (FL) (mcmahan2017communication; kairouz2021advances; jere2020taxonomy) scenario, the situation becomes worse, because training is performed by (potentially many) data providers, rather than a single model provider. Claims about models trained using FL thus require reputational trust in all data providers, as well as the model provider. Nevertheless, FL is an appealing approach in many settings, because it allows data providers to retain control over datasets. Its use has been proposed for applications ranging from training large predictive language models on user input (hard2018federated; ramaswamy2019federated), to training fraud detection models while protecting transaction data of competing financial institutions (yang2019ffd).
@@
In this paper, we explore the question whether it is possible “to provide verifiable claims about how a model was trained using FL by multiple data providers”. We observe that this would require systems support during training to discover any deviations by parties from the agreed training protocol, which would invalidate specific claims. For example, the system would have to discover if a data provider skips a given pre-processing step that removes personal information from a dataset. There could be various reasons for such deviations, ranging from accidental (e.g., bugs in the code) through to actively malicious (e.g., deliberately skipping steps to reduce costs).
@@
A strawman design for creating an FL training system with verifiable claims about the trained FL model would be to execute the FL training computation inside trusted execution environments (TEEs), such as Intel SGX enclaves (arnautov2016scone) or AMD SEV-SNP confidential virtual machines (VMs) (sev2020strengthening). TEEs ensure that the integrity and confidentiality of computation is preserved by the hardware: they isolate the TEE memory used for the computation and offer attestation proofs, which contain a signed measurement of the TEE memory taken at deployment time. Indeed, prior designs of FL training systems have used TEEs to offer claims about trained ML models by leveraging such attestation proofs (openfl; quoc21secfl; guo2024trustworthy).
@@
We argue, however, that TEEs are the wrong abstraction to create verifiable claims about FL model training for two reasons:
@@
(1) Existing solutions focus on static attestation proofs. In FL training systems with TEEs (openfl; quoc21secfl; guo2024trustworthy), TEEs produce memory measurements at deployment time to provide a static attestation proof that establishes that the TEE runs the expected FL system and training code. A model provider can then check each data provider’s static attestation e.g., when establishing an initial TLS network connection (openfl). Such static attestation proofs, however, are hard to relate to claims about the final trained model: verifiable claims depend on how the training process was carried out at runtime but a static attestation proof only capture the state at the start of the training process. For example, if a data provider skips data sanitization in a given training round to poison the model, this would be difficult to discern from a single static attestation proof at deployment time.
(2) Existing attestation relies on confidentiality. As we show, TEE implementations rely on confidentiality to provide unforgeable attestation proofs. TEEs attempt to guarantee confidentiality by isolating TEE memory from external access. Unfortunately, this is challenging due to advanced attacks against TEE confidentiality, e.g., by exploiting side channels (foreshadow; zombieload; li2022systematic). If an attacker compromises the confidentiality of a TEE, they will have also compromised the integrity of its attestation proofs. For example, a distributed FL training system that relies on attested TLS connections between data providers and the model provider (guo2024trustworthy) would be vulnerable to an adversarial data provider who manages to leak the TLS session key from their TEE using a side-channel. This would allow them to deviate from the correct training protocol without being detected because they could forge the static attestation proof.
We observe that FL training should instead rely on a set of runtime attestation proofs about data transformations performed during FL training. The fact that particular data transformations, such as data sanitization (granite), local training (kairouz2021advances), poisoning mitigations (yin2018byzantine; flshield; gabrielli2023protecting) and model aggregation (yin2018byzantine), have been carried out faithfully can then be related to claims about the final trained FL model. Importantly, these runtime attestation proofs only require guarantees about the integrity of data transformations and not their confidentiality. Therefore, our approach can avoid the more challenging security model of TEEs, which must protect against both confidentiality and integrity attacks.
To realise the above approach, our work makes the following two contributions:
(1) We describe a new abstraction called an exclave, which is a hardware-based integrity-only execution environment. The security guarantees of exclaves do not depend on keeping their data secret from the exclave operator. Unlike TEE implementations, the integrity of exclaves is still preserved even if the exclave operator reads all data within the exclave, either directly or through a side channel. Exclaves support attestation through a hardware-based mechanism (e.g., a secure co-processor (sev2020strengthening)) to measure code integrity and sign an attestation statement at runtime that can include other runtime data. The signing keys for attestation are used and maintained by the hardware only, and are never accessible to software, making exclaves resilient to data leakage attacks. We show how exclaves can be realized through minor modifications to existing hardware-based TEE implementations.
@@
(2) Using exclaves, we describe the design and implementation of VerifiableFL, a system for executing FL training jobs that provides verifiable claims about trained FL models. VerifiableFL uses exclaves to generate runtime attestation proofs of all executed data transformations in the FL job, covering the transformations performed by the data providers and the aggregations performed by the model provider in each training round. Each attestation proof, which we call an exclave data record (EDR), binds together through hashes: (i) the data inputs to the transformation; (ii) the code performing the transformation; and (iii) the resulting data output of the transformation. The EDRs are collected and used to construct an end-to-end exclave dataflow graph (EDG) to establish global properties. The EDG provides a complete and integrity-protected representation of all data transformations that led to the final trained FL model. It can be used by a third-party auditor to verify claims about the FL model, e.g., that data pre-processing and model training has been carried out faithfully by data providers across rounds and that the model provider has aggregated all model contributions by the data providers without bias.
We create a prototype implementation of VerifiableFL that extends the existing FL framework, NVFlare (roth2022nvidia), to execute individual FL tasks, which carry out data transformations in an FL training job, inside exclaves. VerifiableFL achieves this without requiring changes to the FL jobs or the framework implementation. Since our exclave design requires minor changes to proprietary TEE firmware, we implement a variant of our exclave design on AMD SEV-SNP using its virtual TPM feature, thus offering similar performance characteristics but slightly weaker security guarantees.
@@
Our experiments show that VerifiableFL can provide verifiable claims such that the FL model was trained with the correct algorithms and datasets with low overhead: the execution of the FL training job within exclaves and the generation of runtime attestation proofs reduces training throughput by less than 12% compared to a baseline without exclaves.
@@ 2. Verifiable Claims in Federated Learning
Next, we provide background on training machine learning models using federated learning (§2.1) and discuss the challenges when providing verifiable claims about trained models (§2.2). After that, we introduce trusted execution environments (§2.3) and survey whether such approaches (or other techniques) can provide verifiable claims about trained models (§2.4). Finally, we consolidate all of the above into a precise threat model and set of security requirements for our desired solution (§2.5).
@@ 2.1. Federated learning
Techniques exist to ensure the protection of datasets or the quality of the trained FL model. For example, data providers may apply differential privacy (DP) (dwork2014algorithmic) to bound the contribution of individual data records in the training dataset. In DP, a data provider adds noise to the model updates to protect the privacy of its dataset. In addition, to ensure the quality of the trained FL model, FL jobs may include algorithmic approaches to detect or mitigate against poisoning attacks (tolpegin2020data), e.g., using outlier detection (tolpegin2020data) or robust aggregation (blanchard2017machine; yin2018byzantine).
2.2. Claims about trained FL models
There is an increasing requirement to provide information about aspects of the training process for ML models. Proposals for providing such information, which are now being used in practice, include datasheets for datasets (datasheetsdatasets) and model cards (modelcard). Analogous to their use in electronics, the idea of datasheets for datasets is to describe the composition and intended use of a training dataset. Model cards perform a similar function, but for a trained model, as shown in the example in Fig. 2. Notably, model cards should ideally include information such as which training algorithm was used, which training datasets were used (e.g., by referring to the datasets’ datasheets), as well as details of any pre-processing performed on the datasets before training (e.g., to filter out certain types of content).
@@ 2.2. Claims about trained FL models
In fully open-source models, the exact training code and dataset are usually made publicly available. In open-weight or closed-source models, the training code and dataset may be confidential, but model providers may still make public claims about the training code, dataset, and data pre-processing. For example, the card for OpenAI’s GPT-5 model claims that their data processing pipeline “includes rigorous filtering to maintain data quality and mitigate potential risks” and that they use “advanced data filtering processes to reduce personal information from training data” (gpt5). For all types of models, information about training may need to be provided to regulators. For example, Article 53 of the EU AI Act requires model providers to provide this information, “upon request, to the AI Office and the national competent authorities” (EUAIAct).
Making claims in FL training, e.g., that particular datasets or attack mitigations have been used, is even more challenging. Claims about a trained FL model require trust in the model provider and all data providers who contributed to the training. There could be various reasons for any of these participants to deviate from the agreed training protocol, either unintentionally or deliberately. For example, a data provider may accidentally omit or misconfigure steps in the pre-processing of the training data, leading to undesirable data being used for training. Alternatively, a data provider may promise to train on a particular dataset (e.g., patient data) but may only train on a subset of this data to reduce costs, resulting in a discrepancy between the claimed and actual training dataset.
It is an open question how to verify such claims about model training, even with access to the training code and datasets. For closed-source models that are only accessible via inference APIs, there is no way to verify claims about the model training; for open-source models, one could attempt to repeat the training process, but depending on the size of the model and the length of training, this may require significant hardware resources and thus incur substantial costs. Even if this were feasible, there is no guarantee that the trained model would match the original model, due to the inherent non-determinism in the training process. Therefore, claims about model training typically have to be accepted based on reputational trust in the entity making the claim. In an FL scenario, however, this reputational trust would have to extend to all participants, i.e., the model provider and also all data providers.
@@ 2.3. Trusted execution environments
Trusted execution environments (TEEs) try to ensure both the confidentiality and integrity of execution in a verifiable manner. They isolate the code and its data, preventing even privileged attackers, such as malicious administrators, operating system (OS) kernels, or hypervisors, from accessing the data or tampering with execution. TEEs also support attestation to enable external parties to verify a measurement taken over the TEE memory, which includes both the code and data. TEEs, including attestation, are implemented using hardware-based mechanisms, such as memory encryption, memory access checks, and hardware signatures for attestation.
TEEs exist at different granularities: ARM TrustZone (trustzone) provides a separate trusted world; Intel Software Guard Extensions (SGX) provides isolation at a process granularity using enclaves. More recently, VM-based TEEs, such as AMD SEV-SNP (sev2020strengthening), Intel TDX (mishra24tdx; cheng2023intel), and ARM CCA (cca24arm; li2023enabling), isolate entire VMs from the hypervisor and other privileged system software.
Hardware support for TEEs has also been integrated into GPUs. For example, NVIDIA’s H100 GPUs support a unified TEE, spanning the CPU-based TEE and the GPU (dhanuskodi2023creating; roth2022nvidia; gpublog). It protects sensitive GPU state, including GPU memory and configuration registers, from external access by the hardware, and all communication between the CPU-based TEE and the GPU is encrypted and integrity-protected. Graphcore IPUs provide similar TEE functionality (vaswani2023confidential).